OpenStudy (anonymous):

directions are simplify each radical if possible if imaginary rewrite in terms of i and simplify? Then it has the square root of -16

OpenStudy (anonymous):

Your question is not clear.

OpenStudy (anonymous):

the answer comes out to 4i so if you have the square root of 27 what would you get

OpenStudy (kinggeorge):

\[\sqrt{-16} = \sqrt{16} \sqrt{-1} = 4i\]

OpenStudy (kinggeorge):

\[\sqrt{27} = \sqrt{3^2 \cdot 3} = 3 \sqrt3\]

OpenStudy (anonymous):

now the square root of 72 and how do you know what numbers to use King George you are doing great!

OpenStudy (anonymous):

??

OpenStudy (kinggeorge):

The key to doing this is to completely factor the radicand. So 72 in this case. To completely factor 72, we first divide by 2, so we have \[72 = 2 \times 36\] We then repeat this process until we it is no longer divisible by 2. \[72 = 2 \times 36 = 2 \times 2 \times 18 = 2 \times 2 \times 2 \times 9 = 2^3 \times 9\] Now we factor 9, so \[72 = 2^3 \times 9 = 2^3 \times 3 \times 3 = 2^3 \times 3^2\] Then we can easily take the square root of this. \[\sqrt{72} = \sqrt{2^3 \times 3^2} = \sqrt{2 \times (2^2 \times 3^2)} = \sqrt2 \times (2 \times 3) = 6\sqrt2\]

OpenStudy (kinggeorge):

If you take the square root of some number with an even exponent, simply divide the exponents by two.

OpenStudy (anonymous):

thank you so much!! You have been a big help!

OpenStudy (kinggeorge):

no problem whatsoever.

OpenStudy (anonymous):

what about a square root of -45/36

OpenStudy (anonymous):

these are hard

OpenStudy (kinggeorge):

\[\sqrt{-{{45} \over {36}}} = {\sqrt{-45} \over \sqrt{36}} = {3i \sqrt5 \over 6} = {i \sqrt5 \over 2}\] I'm skipping some of the steps I've told you about just before, but we know that the root of 36 is 6, and 45 is 5 times 9

OpenStudy (anonymous):

wow thanks again!!!!!!!

OpenStudy (kinggeorge):

you're very welcome

OpenStudy (anonymous):

I hope I can find you on more often!!!

OpenStudy (kinggeorge):

I'm on every now and then. I'm making an effort to be here more often this semester.

OpenStudy (anonymous):

good because a lot of people could benefit from your help! you should be a paid tutor

OpenStudy (kinggeorge):

I might. If I have to get another job, I'll definitely consider it.

OpenStudy (anonymous):

how about -49/75

OpenStudy (anonymous):

I know how to do the first step of putting the square root of -49 /square root of 75 but what then

OpenStudy (anonymous):

I dont understand the i part

OpenStudy (kinggeorge):

well if you factor 49, you get 7 x 7. So \[-49 = {-1} \times 7^2\] so the square root of the whole thing is just\[\sqrt{-49} = 7\sqrt{-1}\] This is where the i comes in. Since \[i = \sqrt{-1}\]\[\sqrt{-49} = 7\sqrt{-1} = 7i\] That's really all i is. It's a pretty simple concept once you understand it.

OpenStudy (anonymous):

you should apply to be an open study ambasador of global internApply to be an OpenStudy Ambassador or Global Intern More information in the study groups. Click here to apply! https://openstudy.wufoo.com/forms/openstudy-global-interns-application/ https://openstudy.wufoo.com/forms/openstudy-ambassador-application/

OpenStudy (anonymous):

so what is the answer for that problem then

OpenStudy (kinggeorge):

since \[\sqrt{-49} = 7i\]and\[\sqrt{75} = \sqrt{25 \times 3} = 5\sqrt3\]\[\sqrt{-49 \over 75} = {7i \over 5\sqrt3} = {7i \sqrt3 \over 15}\]That last step you can get by multiplying by \[\sqrt3 \over \sqrt3\]

OpenStudy (kinggeorge):

I'll be back in a few minutes if you still have more questions.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

That is all for tonight so thank you so much!!! You ought to be a paid helper! Check into that posting I sent you it came up on my screen so you should really do it!